OpenStudy (anonymous):

Help for a medal?!?! The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (anonymous):

f(3)= -3 f(2) -2 f(1) f(3) = ? What did you get?

OpenStudy (anonymous):

f(4) = -3 f(3) -2 f(2) f(4) = ?

OpenStudy (anonymous):

f(2) = 12 f(1) = 4 Plug in

OpenStudy (anonymous):

Sorry, I was eating lunch. I'm not sure, I have had a difficult time trying to answer these questions. I'm confused on how to do these

OpenStudy (anonymous):

@study100 that's right!

OpenStudy (anonymous):

To get f(3) , plug in f(2) and f(1). f(3) = -3 (12) - 2(4) f(3) = ?

OpenStudy (anonymous):

Would it be -28?

OpenStudy (anonymous):

@study100 ?

OpenStudy (anonymous):

@study100 ?!?!

OpenStudy (anonymous):

@study100 ?!?!??!?!?!

OpenStudy (anonymous):

Yes it would, sorry didn't see the notifications

OpenStudy (anonymous):

and now that you have f(3) = -28 use it to find f(4)

OpenStudy (anonymous):

f(4) = -3 f(3) - 2f(2) f(4) = -3(-28) - 2(12)

OpenStudy (anonymous):

Wow, that took a while :O

OpenStudy (anonymous):

I'm not always online : O sorry...?

OpenStudy (anonymous):

Everytime I tagged you you've been online. Anyways would it be 60?

OpenStudy (anonymous):

I'm sorry, -60?

OpenStudy (anonymous):

It shows online status, but not like "online" online

OpenStudy (anonymous):

and what's -3 times -28

OpenStudy (anonymous):

negative x negative = positive

OpenStudy (anonymous):

84

OpenStudy (anonymous):

84 - 24 = ?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Yes :) Answer A !

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcomed! :) Good job as well.

OpenStudy (anonymous):

Do you mind helping with another one?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = 2m + 3?

OpenStudy (anonymous):

@study100 ...

OpenStudy (anonymous):

I'll take a look right away :) I'm here now.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Given: y= f(x) x= 2m +3 find ordered pair (x,y) step 1: plug 2m +3 into x so (2m+3, f(x) step 2: now plug 2m+3 into f(x) so (2m +3, f(2m+3)) conclusion: (2m+3, f(2m+3) )

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcomed :)